OpenStudy (anonymous):

I've never used the quadratic formula before I don't understand the formula... How would I solve for x? 2x^2-5x-4=0

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\\\] can you find those ?

OpenStudy (anonymous):

2,5,4?

hartnn (hartnn):

not actually, you have to account for - *minus* sign also. so, a =2 is correct, what about b,c ?

OpenStudy (anonymous):

-5 and -4?

hartnn (hartnn):

yes, those are correct, now calculate \(b^2-4ac=....?\)

OpenStudy (anonymous):

I'm not sure :/

hartnn (hartnn):

you have a=2, b= -5, c=-4 just plug in values! \(b^2-4ac=(-5)^2-4(2)(-4)=...?\)

OpenStudy (anonymous):

25-8x-4

hartnn (hartnn):

where does 'x' come from ??

hartnn (hartnn):

or is it multiplication sign ?